OpenStudy (anonymous):

Jacob bought seven bottles of water. Each bottle contains 1/2 liter of water. How many liters did he buy?

OpenStudy (anonymous):

3.5

OpenStudy (bahrom7893):

7*(1/2) = 7/2=3.5

OpenStudy (anonymous):

Simple division

OpenStudy (moongazer):

cool pic

OpenStudy (bahrom7893):

ya i said that too.. earlier haha

OpenStudy (saifoo.khan):

three point five